Project

General

Profile

DecamNews

Quirks and Features - Operational issues you should be aware of. Many Solved.

  1. Imager Vacuum Maintained by Ion Pump (4/8/13)
    The gate valve on the imager side of the turbo pump + roughing pump combination is closed, leaving the imager vacuum on only the ion pump.
  2. SISPI Shutdown Problem (3/12/13)
    SISPI was unable to be shut down through the GUI (from the Master Console session, neither the button nor manually typing the shutdown command into the System Controls worked; the page had been refreshed and that didn't fix it either). The shutdownSISPI command also failed with "unable to reach Name Server" errors (the instance name was copy/pasted from the Architect's terminal, so it was not simply a typo in the instance name, and the command was retried). Since the safe shutdown methods failed, and the vnc window didn't like passing a ctrl-C to the terminal, and vsub was already off and the block filter was in, SISPI was shut down by sending a SIGINT to the architect's main process. It was then restarted successfully (the first restart had a Labview error, but the second time worked fine).
    (Editor's Note: This might be on the SISPI News page instead of this one.)
  3. Focal Plane Temperature Excursion (2/15/13)
    The focal plane temperature had a 1/2 deg C "glitch" on evening of 02/13/13. See Elog 54639. Here's a plot with the temperature on the RTDs over 2 days.
  4. Kentools Update (02/03/13)
    From Steve: I have rewritten the critical part of the "center" command in C, and it now can measure these big offsets successfully and much faster than before. If you would like, I will install this on observer2 (and hopefully leave everything else working). Let him know if there are problems.
  5. Dry Gas Purge (12/12/12)
    The LN2 tank is now being tapped faster to provide a dry gas purge
    of the electronics. This is in addition to purge of the C5 lens and
    high pressure dry gas for the filter changer.
  6. Readout Crate Frequent Fan Error Alarms Masked (12/12/12)
    (New) Crate 3 Fan Number 4, (Old) Crate 4 Fan Number 4
    The fans are turning but the fan monitors are not reliable.
  7. Ion Pump Controller Turned On (12/11/12)
    The ion pump on the camera was enabled. Tests are underway as to whether this puts any additional
    RO noise on the CCDs. Early indications is that there's insignificant addtional noise. Late ones, too.
  8. DECal Restart Procedure (11/26/2012)
    DECal stopped after either a network problem or receipt of a bad command. Here is the procedure for starting it back up.
    Restart procedure
  9. CCD/Electronics Problems/Issues/Questions (11/19/2012)
    1. CCD N30: poor CTI, particularly bad behavior. Maybe the problem is connections?
    2. Two of the Focus Chips: one low amplifier on each
    3. Several CCDs: non-linearity at very low light levels. We have an idea tfor a fix that we need to test at FNAL 1st.
    4. Test RO Noise with Ion Pump Controller in new Location
    5. Disable Some RO Crate Fan Monitors
    6. Outstanding question (12/04/12): can we clock the serial registers during an exposure?
  10. Dry gas Purge of Monsoon Crates (12/04/2012)
    The electronics crates are cooled by NESLab Chiller with coolant temperature set to 15C. On previous night we got to a dewpoint of 12C. This difference is smaller than desired. We temporarily increased the coolant temperature to 18C. the plan is to implement a dry gas purge using LN2 boil off ASAP and then to set the coolant temperature back to nominal. We use an administrative limit of 6 deg C between the NESLab chiller supply temperature and the dewpoint. That is - we close up when the dewpoint hits 12 C.
    Actually, TELOPs set the NESLab back to 15 C in the afternoon, so that's where we are at present.
  11. Guider Noise (Sept. 13, 2012)
    Clocking the guide CCDs during 2kx4k RO caused noise in the images. This was resolved by changing the SISPI sequence so the the guide CCDs didn't clock while the image CCDs were reading out.

Files